# Advances in Protein Representation Learning: Methods, Applications, and Future Directions

### v1

@@ Abstract
Enzymes are biological catalysts that accelerate chemical reactions by lowering activation energy, playing fundamental roles in metabolism, signal transduction, and biomolecule synthesis and degradation. Their high specificity and efficiency make them indispensable in both cellular processes and industrial applications, including pharmaceutical production, biofuel synthesis, food processing, and environmental remediation. Enzyme design focuses on engineering or optimizing enzymes to enhance catalytic efficiency, stability, and substrate specificity for targeted reactions. Traditional approaches such as directed evolution and rational design have been widely applied but often require extensive experimental screening 171. Computational strategies now enable in silico enzyme prediction and optimization, facilitating a more efficient design process.
@@
Benchmark datasets play a critical role in assessing enzyme design models by providing standardized evaluation frameworks for sequence-function relationships and catalytic performance. UniProt 102 offers a comprehensive database of protein sequences and functional annotations across diverse organisms, supporting sequence-based enzyme design and optimization. BRENDA 172 compiles detailed enzyme data, including sequence, structure, and kinetic parameters, enabling validation of catalytic efficiency, stability, and substrate specificity in designed enzymes. These benchmarks provide essential resources for the evaluation and development of computational enzyme engineering strategies.
@@
Recent advancements in enzyme design for specific Enzyme Commission (EC) classes have demonstrated promising results, with models generating sequences that closely resemble reference enzymes while maintaining desired catalytic functions 28, 173. Beyond EC-based classification, alternative approaches have explored de novo enzyme generation by assembling active site and scaffold libraries, followed by refinement algorithms to improve functionality 174. While EC classification provides a structured framework, relying solely on predefined categories may limit model generalization to novel, unseen reactions. To address these challenges, recent research has shifted toward reaction-conditioned enzyme design, which directly models enzyme-substrate relationships rather than relying on predefined EC classifications 175, 176. This approach enables greater flexibility in enzyme generation, allowing models to learn catalytic patterns from reaction-specific data.
@@
Another key challenge in enzyme sequence design is the limited understanding of enzyme-substrate catalytic mechanisms. Even when designed enzyme sequences fold correctly into 3D structures, catalytic pocket formation and binding interactions often remain poorly characterized. Recent efforts have integrated generative models for enzyme scaffolds, active sites, and protein language models to improve enzyme-substrate interaction modeling 177. By generating enzyme-substrate binding structures, these methods aim to refine catalytic mechanism prediction and support the design of enzymes capable of catalyzing novel reactions.
8.4 8.4. Drug Discovery
@@
To support model development and evaluation, benchmark datasets provide standardized training and testing resources. CrossDocked 178 consists of cross-docked protein-ligand pairs, challenging models to generalize across diverse binding poses, a crucial ability for designing molecules across multiple protein targets. Binding MOAD (Mother of All Databases) 179 offers a large collection of high-resolution protein-ligand complexes with experimentally measured binding affinities, serving as a foundation for refining docking precision and affinity predictions. Together, these datasets facilitate robust evaluation of structure-based drug design (SBDD) models.
Recent advances in PRL have led to deep learning models that generate molecular structures conditioned on 3D representations of protein binding pockets. These models capture the geometric and physicochemical attributes of binding sites, facilitating the design of small molecules with high affinity and specificity. A predominant strategy in structure-based drug design (SBDD), this pocket-conditioned molecular generation paradigm includes notable models such as Pocket2Mol 180, ResGen 181, PocketFlow 182, DeepICL 183, and DiffSBDD 184, all of which utilize 3D structural data to guide molecular design. By anchoring molecular generation to the spatial and chemical features of protein pockets, these approaches improve ligand design precision and streamline SBDD workflows.
Despite these advances, accurately defining binding pockets for novel protein targets remains a challenge, particularly when structural annotations are unavailable or ambiguous. While many models rely on predefined pockets, emerging efforts are shifting toward whole-protein drug design, allowing models to identify potential binding regions in a data-driven manner without requiring prior annotations 72. This approach broadens the applicability of SBDD to less-characterized targets. Additionally, a complementary research direction focuses on sequence-based drug design, where molecular structures are generated directly from protein sequences, bypassing the need for 3D structural data 185, 186. These methods leverage Protein Language Models (PLMs) to extract sequence-level functional insights, offering an alternative strategy for designing drugs against targets with limited or unreliable structural information.
8.4.2 8.4.2. Drug Repurposing
@@
To facilitate standardized evaluation, benchmark datasets provide essential validation resources for both virtual screening and DTI-based approaches. For virtual screening, DUD-E (Directory of Useful Decoys: Enhanced) 187 and LIT-PCBA 188 are widely used. DUD-E pairs protein targets with both active compounds and decoys, assessing a model’s ability to differentiate true binders from non-binders in complex screening tasks. LIT-PCBA, derived from PubChem bioassays, offers experimentally validated active and inactive compounds, reflecting the challenges encountered in real-world drug repurposing efforts. For DTI-based link prediction, datasets such as PrimeKG 189, Therapeutics Data Commons (TDC) 190, BindingDB 191, and BioSNAP 192 provide extensive collections of validated drug-target interactions. These resources support the development and benchmarking of computational models that predict novel drug-target relationships, enabling systematic drug repurposing.
@@
Virtual screening can be further categorized into docking-based and similarity-based methods. Docking-based approaches use molecular docking simulations, such as AutoDock Vina 193, to computationally dock large drug libraries against protein targets and predict binding affinities. While effective, traditional docking is computationally intensive and limits scalability for high-throughput screening. To overcome these limitations, deep learning-based docking models, including P2Rank 194, EquiBind 195, TANKBind 196, DiffDock 197 and HelixDock 198, have been developed to predict binding poses and affinities with reduced computational costs. These models leverage 3D protein-ligand complexes to learn structural binding interactions and improve docking accuracy. In contrast, similarity-based approaches rely on the “similarity principle”, which assumes that structurally similar molecules exhibit similar biological properties 199. Rather than explicitly docking compounds, these methods use learned molecular representations to rapidly identify structurally related drug candidates. Recent advances in PRL have enabled drug and protein embeddings to be mapped into a shared representation space, facilitating efficient similarity searches. Unlike docking-based methods that depend on 3D structural data, similarity-based approaches do not require explicit protein-ligand modeling. Instead, they employ pretrained models, such as Protein Language Models (PLMs) and Molecular Graph Neural Networks (GNNs), to encode proteins and ligands separately. Notable examples include ConPLex 200, DrugCLIP 201, and SPRINT 202, which use contrastive learning to map proteins and drugs into a joint representation space, allowing for rapid and scalable virtual screening without the computational cost of docking simulations.
@@
Beyond virtual screening, DTI-based link prediction offers an alternative paradigm for drug repurposing by formulating drug-target interactions as a graph-based link prediction task. In this approach, drugs and proteins are represented as nodes, while known interactions form edges. A key challenge in DTI-based models is how drug and protein nodes are represented, as the quality of node embeddings directly impacts model performance. Recent methods have introduced various node embedding strategies, which can be classified into sequence-based and multimodal embeddings. Sequence-based embeddings use PLMs for protein sequences and SMILES-based transformers for molecular representations 203, 204, encoding drugs and proteins based on their primary sequences without requiring structural data. Meanwhile, multimodal embeddings integrate diverse biological information sources, combining sequence, structure, and auxiliary biological data, such as functional annotations, side-effect similarities, and pathway interactions, to construct more comprehensive representations 205, 206. By capturing both structural and functional relationships, these models improve the predictive accuracy of link prediction tasks, making DTI-based drug repurposing a valuable complement to virtual screening approaches by providing insights into drug-target associations that may not be captured through traditional binding-based models.
9 9. Future Directions and Open Challenges
@@
9.1 9.1. Biological Sequence Representation Learning
@@
The representation learning of DNA and RNA shares foundational similarities with protein representation learning, as both involve extracting meaningful features from biological sequences to predict function, interactions, and structural properties. Like proteins, DNA and RNA sequences follow a linear arrangement of biological “letters” that encode essential biological information. Advances in self-supervised learning and transformer-based architectures, which have successfully modeled protein sequences, can similarly be applied to DNA and RNA to capture sequence patterns, evolutionary conservation, and functional motifs. Recent efforts to develop genome language models have shown promising results, demonstrating their potential in learning meaningful representations of genomic sequences and improving the prediction of gene function, regulatory elements, and chromatin accessibility 207, 208, 209.
However, DNA and RNA present unique challenges distinct from proteins. Unlike proteins, which have a well-defined alphabet of 20 amino acids and a structured hierarchical organization, DNA and RNA sequences consist of only four nucleotides (A, T/U, G, and C) but often span much longer sequences, reaching tens of thousands or even millions of base pairs. This length disparity makes learning effective representations computationally intensive and necessitates efficient sequence compression strategies. Additionally, while proteins fold into stable three-dimensional structures that largely dictate function, DNA and RNA often exhibit dynamic secondary and tertiary structures influenced by sequence context, environmental factors, and cellular conditions. Capturing these structural variations requires novel modeling approaches that integrate both sequence-based and structural information. Furthermore, the function of DNA and RNA is heavily modulated by epigenetic modifications, chromatin accessibility, and interactions with regulatory proteins. Unlike proteins, where function is often inferred from sequence and structure, the functional state of a genomic region is highly context-dependent, requiring models to account for regulatory mechanisms beyond the primary sequence.
9.2 9.2. Scaling PRL Models for Large-Scale Applications
@@
One promising direction for improving scalability is model distillation, a technique that has seen significant progress in Natural Language Processing (NLP) and Computer Vision (CV). In NLP, approaches such as DistilBERT 210 and TinyBERT 211 have successfully compressed large transformer models while preserving much of their predictive power. Similarly, in CV, TinyViT 212 and DearKD 213 leverage distillation to create more compact and computationally efficient models. These techniques demonstrate the potential of distilling knowledge from large, computationally expensive models into smaller, more efficient variants without significant loss of accuracy. Recent efforts have explored applying model distillation techniques to Protein Representation Learning (PRL) to improve scalability and efficiency. For instance, OpenFold 128 employs knowledge distillation to create a more computationally efficient version of AlphaFold.
@@
Applying model distillation to PRL could enhance efficiency by enabling smaller models to retain essential biological insights while significantly reducing memory and computational overhead. This would be particularly beneficial for large-scale virtual screening, protein design, and low-data scenarios where fine-tuning massive models is impractical. Additionally, integrating distillation with techniques such as low-rank adaptations, mixed-precision training, and distributed learning could further enhance efficiency while maintaining high performance.
9.3 9.3. Improving Generalization of PRL Models
@@
While scalability focuses on computational efficiency, generalization remains a critical challenge in PRL. Many models struggle to generalize to unseen proteins, such as those from novel species, synthetic biology applications, or de novo protein designs, limiting their applicability in biomedical and biotechnological research. Beyond handling completely novel proteins, PRL models must also account for mutated proteins, where small amino acid substitutions, deletions, or insertions can significantly alter a protein’s function, stability, or interactions. Understanding how mutations affect learned representations is crucial for applications such as disease mutation analysis, where genetic variants influence protein function; protein engineering, where specific mutations can optimize protein stability and activity; and drug resistance prediction, where mutations in pathogens impact therapeutic efficacy. However, the availability of labeled data for both novel proteins and mutant proteins is often extremely limited, making it difficult to train models that can reliably predict their properties.
@@
To address this, zero-shot and few-shot learning approaches need to be developed, enabling PRL models to generalize effectively to unseen proteins and mutations with minimal supervision. Recent studies have demonstrated promising results in this area, showing that pretrained PRL models can infer functional properties and structural effects of mutations even in data-scarce settings, improving predictive performance in mutation effect prediction and de novo protein engineering 214, 57.
9.4 9.4. Enhancing Explainability in PRL Models
@@
Despite their success, many state-of-the-art PRL models operate as black boxes, making it difficult to interpret how they learn protein representations and make predictions. This lack of transparency hinders scientific discovery, limits trust in model outputs, and reduces their applicability in critical fields such as medicine, biotechnology, and drug development. As PRL models become increasingly complex—leveraging deep neural networks, large-scale pretraining, and self-supervised learning—there is a growing need for approaches that can make their predictions more interpretable and biologically meaningful.
@@
One approach to enhancing explainability in PRL is the development of self-interpretable architectures that inherently capture biologically meaningful representations. Explainable AI techniques, such as attention mechanisms, saliency mapping, and feature attribution, allow models to highlight critical sequence motifs, structural domains, and evolutionary patterns that drive their predictions. By identifying key residues, structural elements, and physicochemical properties relevant to specific functions or interactions, these methods improve model transparency and reliability. This enhanced interpretability facilitates applications in functional protein annotation, biomarker discovery, and rational protein design. Recent studies have explored these approaches, applying various Explainable AI techniques to improve the interpretability of PRL models 215, 216.
@@
Another promising direction is leveraging advances in Natural Language Processing (NLP) and Large Language Models (LLM) to make PRL models more accessible and interpretable. Inspired by recent successes in LLM, PRL models could be designed to generate natural language descriptions of protein functions, structural properties, and interactions, making predictions more intuitive for researchers in biology, medicine, and drug development. Beyond simple text-based outputs, reasoning-based approaches could further enhance explainability by allowing models to logically infer protein functions and molecular interactions rather than relying solely on statistical pattern recognition. Techniques such as retrieval-augmented generation (RAG) and biomedical knowledge graphs (BKGs) could serve as foundational tools for reasoning, enabling models to retrieve relevant biological context from structured databases and scientific literature. Instead of merely generating descriptions, PRL models could use retrieved biological relationships and prior knowledge to construct logical justifications for their predictions, improving their transparency and reliability. Recent studies have already begun exploring these approaches, demonstrating promising results in enhancing model interpretability and providing biologically meaningful insights 217, 218. By integrating LLM-based reasoning, knowledge retrieval, and structured biological insights, PRL models could evolve into more explainable, evidence-driven systems, facilitating their adoption in research and clinical applications.
@@
Figure 5: Key future directions in Protein Representation Learning (PRL). The figure outlines four critical challenges and potential advancements: (i) Expanding PRL to DNA and RNA representation learning to leverage shared methodologies while addressing unique challenges, (ii) Enhancing scalability to improve computational efficiency and accessibility of large-scale models, (iii) Strengthening generalization to ensure robustness across unseen proteins and genetic variations, and (iv) Advancing explainability to improve model interpretability and facilitate trust in biological and biomedical applications.
10 10. Conclusion

### v2

@@ 8.3.4. Enzyme Design
Enzymes are biological catalysts that accelerate chemical reactions by reducing the activation energy and play a fundamental role in metabolism, signal transduction, and biomolecule synthesis and degradation. Their high specificity and efficiency make them indispensable in both cellular processes and industrial applications, including pharmaceutical production, biofuel synthesis, food processing, and environmental remediation. Enzyme design focuses on the engineering or optimization of enzymes to enhance catalytic efficiency, stability, and substrate specificity for targeted reactions. Traditional approaches such as directed evolution and rational design have been widely applied, but often require extensive experimental screening (doi:10.1021/acscentsci.3c01275, ). Computational strategies now enable in silico enzyme prediction and optimization, facilitating a more efficient design process.
@@
Benchmark datasets play a critical role in assessing enzyme design models by providing standardized evaluation frameworks for sequence-function relationships and catalytic performance. UniProt (boutet2007uniprotkb, ) offers a comprehensive database of protein sequences and functional annotations in various organisms, supporting the design and optimization of sequence-based enzymes. BRENDA (chang2021brenda, ) compiles detailed enzyme data, including sequence, structure, and kinetic parameters, enabling validation of catalytic efficiency, stability, and substrate specificity in designed enzymes. These benchmarks provide essential resources for the evaluation and development of computational enzyme engineering strategies.
@@
Recent advancements in enzyme design for specific Enzyme Commission (EC) classes have demonstrated promising results, with models generating sequences that closely resemble reference enzymes while maintaining desired catalytic functions (munsamy2022zymctrl, ; yang2024conditional, ). Beyond EC-based classification, alternative approaches have explored de novo enzyme generation by assembling active site and scaffold libraries, followed by refinement algorithms to improve functionality (hossack2023building, ). While EC classification provides a structured framework, relying solely on predefined categories may limit model generalization to novel, unseen reactions. To address these challenges, recent research has shifted towards reaction-conditioned enzyme design, which directly models enzyme-substrate relationships rather than relying on predefined EC classifications (mikhael2024clipzyme, ; hua2025reactzyme, ). This approach enables greater flexibility in enzyme generation, allowing models to learn catalytic patterns from reaction-specific data.
@@
Another key challenge in enzyme sequence design is the limited understanding of enzyme-substrate catalytic mechanisms. Even when designed enzyme sequences fold correctly into 3D structures, catalytic pocket formation and binding interactions often remain poorly characterized. Recent efforts have integrated generative models for enzyme scaffolds, active sites, and protein language models to improve the modeling of enzyme-substrate interaction (hua2024reaction, ). By generating enzyme-substrate binding structures, these methods aim to refine the prediction of the catalytic mechanism and support the design of enzymes capable of catalyzing novel reactions.
8.4. Drug Discovery
@@ 8.4.1. Drug Design
To support model development and evaluation, benchmark datasets provide standardized training and testing resources. CrossDocked (francoeur2020three, ) consists of cross-docked protein-ligand pairs, challenging models to generalize across diverse binding poses, a crucial ability for designing molecules across multiple protein targets. Binding MOAD (Mother of All Databases) (benson2007binding, ) offers a large collection of high-resolution protein-ligand complexes with experimentally measured binding affinities, serving as a foundation for refining docking precision and affinity predictions. Together, these datasets facilitate robust evaluation of structure-based drug design (SBDD) models.
Recent advances in PRL have led to deep learning models that generate molecular structures conditioned on 3D representations of protein binding pockets. These models capture the geometric and physicochemical attributes of the binding sites, facilitating the design of small molecules with high affinity and specificity. A predominant strategy in structure-based drug design (SBDD), this pocket-conditioned molecular generation paradigm includes notable models such as Pocket2Mol (peng2022pocket2mol, ), ResGen (zhang2023resgen, ), PocketFlow (jiang2024pocketflow, ), DeepICL (zhung20243d, ), and DiffSBDD (schneuing2024structure, ), all of which utilize 3D structural data to guide molecular design. By anchoring molecular generation to the spatial and chemical features of protein pockets, these approaches improve ligand design precision and streamline SBDD workflows.
Despite these advances, accurately defining binding pockets for novel protein targets remains a challenge, particularly when structural annotations are unavailable or ambiguous. While many models rely on predefined pockets, emerging efforts are shifting toward whole-protein drug design, allowing models to identify potential binding regions in a data-driven manner without requiring prior annotations (ngo2024multimodal, ). This approach broadens the applicability of SBDD to less-characterized targets. Furthermore, a complementary research direction focuses on sequence-based drug design, where molecular structures are generated directly from protein sequences, bypassing the need for 3D structural data (chen2023sequence, ; creanza2025transformer, ). These methods leverage Protein Language Models (PLMs) to extract sequence-level functional insights, offering an alternative strategy for designing drugs against targets with limited or unreliable structural information.
8.4.2. Drug Repurposing
@@
To facilitate standardized evaluation, benchmark datasets provide essential validation resources for both virtual screening and DTI-based approaches. For virtual screening, DUD-E (Directory of Useful Decoys: Enhanced) (mysinger2012directory, ) and LIT-PCBA (tran2020lit, ) are widely used. DUD-E pairs protein targets with both active compounds and decoys, evaluating a model’s ability to differentiate true binders from non-binders in complex screening tasks. LIT-PCBA, derived from PubChem bioassays, offers experimentally validated active and inactive compounds, reflecting the challenges encountered in real-world drug repurposing efforts. For DTI-based link prediction, datasets such as PrimeKG (chandak2023building, ), Therapeutics Data Commons (TDC) (huang2021therapeutics, ), BindingDB (zagidullin2019drugcomb, ), and BioSNAP (biosnapnets, ) provide extensive collections of validated drug-target interactions. These resources support the development and benchmarking of computational models that predict novel drug-target relationships, enabling systematic drug repurposing.
@@
Virtual screening can be further categorized into docking-based and similarity-based methods. Docking-based approaches use molecular docking simulations, such as AutoDock Vina (trott2010autodock, ), to computationally dock large drug libraries against protein targets and predict binding affinities. While effective, traditional docking is computationally intensive and limits scalability for high-throughput screening. To overcome these limitations, deep learning-based docking models have been developed, including P2Rank (krivak2018p2rank, ), EquiBind (stark2022equibind, ), TANKBind (lu2022tankbind, ), DiffDock (corso2023diffdock, ) and HelixDock (liu2023pre, ), to predict binding poses and affinities with reduced computational costs. These models leverage 3D protein-ligand complexes to learn structural binding interactions and improve docking accuracy. In contrast, similarity-based approaches rely on the “similarity principle”, which assumes that structurally similar molecules exhibit similar biological properties (bender2009similar, ). Rather than explicitly docking compounds, these methods use learned molecular representations to rapidly identify structurally related drug candidates. Recent advances in PRL have enabled drug and protein embeddings to be mapped into a shared representation space, facilitating efficient similarity searches. Unlike docking-based methods that depend on 3D structural data, similarity-based approaches do not require explicit protein-ligand modeling. Instead, they employ pre-trained models, such as Protein Language Models (PLMs) and Molecular Graph Neural Networks (GNNs), to encode proteins and ligands separately. Notable examples include ConPLex (singh2023contrastive, ), DrugCLIP (gao2023drugclip, ), and SPRINT (mcnutt2024sprint, ), which use contrastive learning to map proteins and drugs into a joint representation space, allowing for rapid and scalable virtual screening without the computational cost of docking simulations.
@@
Beyond virtual screening, DTI-based link prediction offers an alternative paradigm for drug repurposing by formulating drug-target interactions as a graph-based link prediction task. In this approach, drugs and proteins are represented as nodes, while known interactions form edges. A key challenge in DTI-based models is how the drug and protein nodes are represented, as the quality of the node embeddings directly impacts model performance. Recent methods have introduced various node embedding strategies, which can be classified into sequence-based and multimodal embeddings. Sequence-based embeddings use PLMs for protein sequences and SMILES-based transformers for molecular representations (daza2023bioblp, ; djeddi2023advancing, ), encoding drugs and proteins based on their primary sequences without requiring structural data. Meanwhile, multimodal embeddings integrate diverse sources of biological information, combining sequence, structure, and auxiliary biological data, such as functional annotations, side-effect similarities, and pathway interactions, to construct more comprehensive representations (dang2025multimodal, ; dong2023multi, ). By capturing both structural and functional relationships, these models improve the predictive accuracy of link prediction tasks, making DTI-based drug repurposing a valuable complement to virtual screening approaches by providing insights into drug-target associations that may not be captured through traditional binding-based models.
9. Future Directions and Open Challenges
@@
9.1. Biological Sequence Representation Learning
@@
The representation learning of DNA and RNA shares foundational similarities with the protein representation learning, as both involve extracting meaningful features from biological sequences to predict function, interactions, and structural properties. Like proteins, DNA and RNA sequences follow a linear arrangement of biological “letters” that encode essential biological information. Advances in self-supervised learning and transformer-based architectures, which have successfully modeled protein sequences, can be applied similarly to DNA and RNA to capture sequence patterns, evolutionary conservation, and functional motifs. Recent efforts to develop genome language models have shown promising results, demonstrating their potential in learning meaningful representations of genomic sequences and improving the prediction of gene function, regulatory elements, and chromatin accessibility (nguyen2024sequence, ; brixi2025genome, ; Consens2025, ).
However, DNA and RNA present unique challenges distinct from proteins. Unlike proteins, which have a well-defined alphabet of 20 amino acids and a structured hierarchical organization, DNA and RNA sequences consist of only four nucleotides (A, T/U, G, and C) but often span much longer sequences, reaching tens of thousands or even millions of base pairs. This length disparity makes learning effective representations computationally intensive and necessitates efficient sequence-compression strategies. Furthermore, while proteins fold into stable three-dimensional structures that largely dictate function, DNA and RNA often exhibit dynamic secondary and tertiary structures influenced by sequence context, environmental factors, and cellular conditions. Capturing these structural variations requires novel modeling approaches that integrate both sequence-based and structural information. Furthermore, the function of DNA and RNA is strongly modulated by epigenetic modifications, chromatin accessibility, and interactions with regulatory proteins. Unlike proteins, where function is often inferred from sequence and structure, the functional state of a genomic region is highly context-dependent, requiring models to account for regulatory mechanisms beyond the primary sequence.
9.2. Scaling PRL Models for Large-Scale Applications
@@
One promising direction for improving scalability is model distillation, a technique that has seen significant progress in Natural Language Processing (NLP) and Computer Vision (CV). In NLP, approaches such as DistilBERT (sanh2019distilbert, ) and TinyBERT (jiao2019tinybert, ) have successfully compressed large transformer models while preserving much of their predictive power. Similarly, in CV, TinyViT (wu2022tinyvit, ) and DearKD (chen2022dearkd, ) leverage distillation to create more compact and computationally efficient models. These techniques demonstrate the potential of distilling knowledge from large, computationally expensive models into smaller, more efficient variants without a significant loss of accuracy. Recent efforts have explored applying model distillation techniques to Protein Representation Learning (PRL) to improve scalability and efficiency. For instance, OpenFold (ahdritz2024openfold, ) employs knowledge distillation to create a more computationally efficient version of AlphaFold.
@@
Applying model distillation to PRL could enhance efficiency by enabling smaller models to retain essential biological insights while significantly reducing memory and computational overhead. This would be particularly beneficial for large-scale virtual screening, protein design, and low-data scenarios where fine-tuning massive models is impractical. Furthermore, integrating distillation with techniques such as low-rank adaptations, mixed-precision training, and distributed learning could further enhance efficiency while maintaining high performance.
9.3. Improving Generalization of PRL Models
@@
While scalability focuses on computational efficiency, generalization remains a critical challenge in PRL. Many models struggle to generalize to unseen proteins, such as those from novel species, synthetic biology applications, or de novo protein designs, limiting their applicability in biomedical and biotechnological research. Beyond handling completely novel proteins, PRL models must also account for mutated proteins, where small substitutions, deletions, or insertions of amino acids can significantly alter the function, stability, or interactions of a protein. Understanding how mutations affect learned representations is crucial for applications such as disease mutation analysis, where genetic variants influence protein function; protein engineering, where specific mutations can optimize protein stability and activity; and drug resistance prediction, where mutations in pathogens impact therapeutic efficacy. However, the availability of labeled data for both novel proteins and mutant proteins is often extremely limited, making it difficult to train models that can reliably predict their properties.
@@
To address this, zero-shot and few-shot learning approaches need to be developed, enabling PRL models to generalize effectively to unseen proteins and mutations with minimal supervision. Recent studies have demonstrated promising results in this area, showing that pretrained PRL models can infer functional properties and structural effects of mutations even in data-scarce settings, improving predictive performance in mutation effect prediction and de novo protein engineering (tan2024retrieval, ; cheng2024zero, ).
9.4. Enhancing Explainability in PRL Models
@@
Despite their success, many state-of-the-art PRL models operate as black boxes, making it difficult to interpret how they learn protein representations and make predictions. This lack of transparency hinders scientific discovery, limits trust in model outputs, and reduces their applicability in critical fields such as medicine, biotechnology, and drug development. As PRL models become increasingly complex, using deep neural networks, large-scale pre-training, and self-supervised learning, there is a growing need for approaches that can make their predictions more interpretable and biologically meaningful.
@@
An approach to enhancing explainability in PRL is the development of self-interpretable architectures that inherently capture biologically meaningful representations. Explainable AI techniques, such as attention mechanisms, saliency mapping, and feature attribution, allow models to highlight critical sequence motifs, structural domains, and evolutionary patterns that drive their predictions. By identifying key residues, structural elements, and physicochemical properties relevant to specific functions or interactions, these methods improve model transparency and reliability. This enhanced interpretability facilitates applications in functional protein annotation, biomarker discovery, and rational protein design. Recent studies have explored these approaches, applying various Explainable AI techniques to improve the interpretability of PRL models (gong2025multigrandti, ; medina2024interpretable, ).
@@
Another promising direction is to leverage advances in Natural Language Processing (NLP) and Large Language Models (LLM) to make PRL models more accessible and interpretable. Inspired by recent successes in LLM, PRL models could be designed to generate natural language descriptions of protein functions, structural properties, and interactions, making predictions more intuitive for researchers in biology, medicine, and drug development. Beyond simple text-based outputs, reasoning-based approaches could further enhance explainability by allowing models to logically infer protein functions and molecular interactions rather than relying solely on statistical pattern recognition. Techniques such as retrieval-augmented generation (RAG) and biomedical knowledge graphs (BKGs) could serve as foundational tools for reasoning, enabling models to retrieve relevant biological context from structured databases and scientific literature. Instead of merely generating descriptions, PRL models could use retrieved biological relationships and prior knowledge to construct logical justifications for their predictions, improving their transparency and reliability. Recent studies have already begun exploring these approaches, demonstrating promising results in improving the interpretability of the model and providing biologically meaningful insights (ma2025prottexstructureincontextreasoningediting, ; wang2025prot2chatproteinllmearly, ). By integrating LLM-based reasoning, knowledge retrieval, and structured biological insights, PRL models could evolve into more explainable, evidence-driven systems, facilitating their adoption in research and clinical applications.
@@
Figure 5. Key future directions in Protein Representation Learning (PRL). The figure outlines four critical challenges and potential advancements: (i) Expanding PRL to DNA and RNA representation learning to leverage shared methodologies while addressing unique challenges, (ii) Enhancing scalability to improve computational efficiency and accessibility of large-scale models, (iii) Strengthening generalization to ensure robustness across unseen proteins and genetic variations, and (iv) Advancing explainability to improve model interpretability and facilitate trust in biological and biomedical applications.
10. Conclusion